OSG Technology Area Meeting, 16 October 2017¶
Coordinates: Conference: 719-284-5267, PIN: 57363; https://www.uberconference.com/osgblin
Attending: BrianL, Derek, Edgar, Mat, Suchandra, TimC, TimT
LIGO announced neutron star collision detection!
- This week: Derek
- Next week: Edgar
- 9 (-3) open tickets
|# of tickets||Δ||State|
|4||+4||Ready for Testing|
|0||-17||Ready for Release|
|Name||Version||Development Freeze||Package Freeze||Release||Notes|
|November||3.4.5, 3.3.30||2017-10-30||2017-11-06||2017-11-14||5 week cycle|
Notes: Additional “urgent” releases may be scheduled for the 4th Tuesday of each month. The Testing date is when acceptance testing will be scheduled for releasable packages; if a package is added after this date, it may not be possible to schedule adequate testing time, thereby forcing it into the next release.
OSG Software Team¶
- Documentation migration remaining items
- SoftwareTeam migrations
- Release3 and SoftwareTeam archivals
- TWiki migration headers and archive doc deletions
- Globus EOL: Actively discussing repo destination with EGI and WLCG
- November release development freeze - 11/30
- Per-page migration headers are preferred but a per-web migration would be easier
- We need to search and replace twiki references in software
- Colorado (BrianL) - BeStMan not accepting RFC-compliant proxies
- JINR (BrianL) - blahp failures due to some issue with ASCII chars
- LBNL (BrianL) - issues with the blahp may have been due to misconfiguration
- Taiwan (BrianL) - idle jobs on CE, likely a basic configuration issue
- GLOW (Derek, Edgar) - GLOW isn't running on many OSG sites we would expect. Edgar is leading the investigation.
- CMS / NERSC (Derek) - Add userspace StashCache for CMS data caching. Ongoing.
OSG Release Team¶
- Tim Theisen will handle the next release if the new procedure is used
|0||+0||4||+4||0||+0||4||+4||Ready for Testing|
|0||+0||0||+0||0||+0||0||+0||Ready for Release|
- Update osg-system profiler
- CA Certificate update handled in 2 days
OSG Investigations Team¶
- Worked on cleaning up GRACC data.
- Major update of GRACC-APEL reports to reflect non-zero CpuDuration in WLCG reports
- Worked on getting data from GRACC for press release (requested by BrianB/FKW)
- More work on writable stashcache.
- Installed CMS data cache at NERSC in userspace. Will cache CMS data destined at NERSC.
- Lots of CVMFS changes on OSG hosted repos (except oasis).
- Finalizing GRACC changes
- Draft GRACC SLA proposal
- Writable StashCache. Enabling access for initial testers.
- Improve CMS cache at NERSC, productionize
None this week